# Consistent Validation for Predictive Methods in Spatial Settings

### v1

@@ Abstract
Spatial prediction tasks are key to weather forecasting, studying air pollution, and other scientific endeavors. Determining how much to trust predictions made by statistical or physical methods is essential for the credibility of scientific conclusions. Unfortunately, classical approaches for validation fail to handle mismatch between locations available for validation and (test) locations where we want to make predictions. This mismatch is often not an instance of covariate shift (as commonly formalized) because the validation and test locations are fixed (e.g., on a grid or at select points) rather than i.i.d. from two distributions. In the present work, we formalize a check on validation methods: that they become arbitrarily accurate as validation data becomes arbitrarily dense. We show that classical and covariate-shift methods can fail this check. We instead propose a method that builds from existing ideas in the covariate-shift literature, but adapts them to the validation data at hand. We prove that our proposal passes our check. And we demonstrate its advantages empirically on simulated and real data.
@@ 1 Introduction
Researchers are often interested in making predictions in a spatial setting. For instance, scientists would like to predict sea surface temperature (SST) for weather forecasting and climate research (Minnett, 2010), predict air pollution at population centers to better understand the effect of pollution on health outcomes such as kidney disease (Remigio et al., 2022), or predict the prevalence of an invasive species for ecological management (Barbet-Massin et al., 2018). Characterizing the reliability of these predictions is key to understanding their suitability for certain downstream applications; e.g., Minnett (2010) describe acceptable SST error tolerances for weather forecasting. Estimating prediction accuracy can also be used to decide among a collection of predictive methods, as in Shabani et al. (2016).
@@
Importantly, in the spatial setting, predictive methods need not always arise from a statistical or machine learning approach built using training data. The predictive method is often a complex physical model provided by a third party (Remigio et al., 2022; Minnett, 2010; Gupta et al., 2018). Or it could combine physics and data-driven models (Banzon et al., 2016; Werner et al., 2019; Özkaynak et al., 2013).
@@
In any of these cases, it is common to estimate the performance of a predictive method by using a set of validation data. More precisely, we are ultimately interested in predicting a response at what we will call test sites; in the SST example above (Minnett, 2010), the test sites are points on a grid (often called a map), or in the air pollution example (Remigio et al., 2022), the test sites are 28 counties in the US Northeast. We can make predictions at the test data, but we do not have access to direct observations of the responses there. We do have observed responses in the validation data, which we assume were not used in forming the predictive method being evaluated; in the SST example, scientists have SST observations taken by boats and buoys as validation data. So our aim is to estimate the average loss (i.e. risk) at the testing data using the validation data. We will see that many popular or natural approaches fail at this task.
@@
One widely used approach, called the holdout, [^1] The name originally referred to “holding out” data for validation, with the remainder of available data going toward training. While we maintain the naming convention, we emphasize that in the present development there need not be any training data. Minnett (2010); Gupta et al. (2018); Duan et al. (2019), among many others, use this approach. estimates the test risk by taking the empirical average of the validation loss. When the validation and test data are independent and identically distributed (i.i.d.) from the same distribution, the holdout has a rigorous justification (Devroye, 1976; Langford, 2005). But in spatial problems, the validation and test sites need not be similarly dispersed, all data may be spatially correlated, and the test sites are often fixed rather than random; recall the grid or point prediction examples above. Indeed, Roberts et al. (2017) have observed problems with the holdout in practice. In the special case where the predictive method is data-driven, some authors (e.g. Telford & Birks, 2005) have suggested choosing holdout validation sites far from the training data sites. But we expect this proposal to still suffer from the problems just mentioned, and indeed simulation studies suggest it can misestimate test risk (Ploton et al., 2020; De Bruin et al., 2022).
@@
In what follows, we start by laying out a precise formulation of the prediction validation task in the spatial setting (Section 2). We formalize a desirable property for test-risk estimators: that, if arbitrarily dense validation data accrues in a region including the test points, the test-risk estimate should become arbitrarily accurate (Section 3). We prove that both the holdout estimator and an estimator advocated in the covariate-shift literature (Loog, 2012; Portier et al., 2023) fail to satisfy this spatial consistency property (Section 4). We propose instead to build on the $k$-nearest neighbor estimator of Loog (2012). In particular, Loog (2012); Portier et al. (2023) advocated fixing $k=1$ for covariate-shift problems. We instead derive an upper bound on the error of the general-$k$ estimator for estimating test risk (Section 5.1); crucially, our bound is conditional on the test and validation sites. We prove that choosing $k$ by optimizing our upper bound yields an estimator with our desired spatial consistency property (Section 5.2). Unlike covariate-shift results (e.g. Portier et al., 2023), our results are directly applicable to problems where the test sites are most reasonably thought of as fixed. We illustrate the accuracy and practicality of our proposed method in simulated and real data analyses (Section 6), including tasks in both grid and point prediction. We discuss further related work in App. A.
@@ 2 Estimating Test Risk in a Spatial Problem
We now formalize risk predictions at test points in a spatial setting. We assume each data point occurs at a spatial location $S\in\mathcal{S}$, where the spatial domain $(\mathcal{S},d_{\mathcal{S}})$ is a metric space. Each data point has observed covariates $X\in\mathcal{X}$ and a response $Y\in\mathcal{Y}$. We treat the covariates as a fixed spatial field, $\chi:\mathcal{S}\to\mathcal{X}$; i.e., the covariates at a point are exactly specified by applying $\chi$ at the spatial location of the point.
@@ 2.1 Test risk of a spatial predictive method
We assume we have access to a predictive method $h:(\mathcal{S},\mathcal{X})\to\mathcal{Y}$. Since we assume covariates are determined by their spatial location, we will find it convenient to define $h^{\chi}:\mathcal{S}\to\mathcal{Y}$ to be the prediction made by $h$ at the location $S$: $h^{\chi}(S)=h(S,\chi(S))$. We suppose that practitioners would like to use $h$ to predict the response at a set of test sites where the response is currently unknown. We will find it useful to collect the $M^{\textup{{test}}}$ test data points, including true but unobserved responses, in $D^{\textup{{test}}}=(S^{\textup{{test}}}_{m},X^{\textup{{test}}}_{m},Y^{\textup{{test}}}_{m})_{m=1}^{M^{\textup{{test}}}}$.
To quantify performance of a predictive method, we need a loss. We will assume the loss is bounded, as is typically the case for practically-bounded responses; cf. temperature, pressure, or other physical quantities. [^2] Loss is also bounded for classification error or robust regression cases such as trimmed squared loss or Tukey’s biweight loss.
@@
The loss $\ell$ is a non-negative, $\Delta$-bounded function. That is $\ell:\mathcal{Y}\times\mathcal{Y}\to[0,\Delta]$.
@@
Let $j=\text{test}$. Let $\chi:\mathcal{S}\to\mathcal{X}$ be a fixed function. For $1\leq m\leq M^{j}$, $X^{j}_{m}=\chi(S^{j}_{m})$ and $Y_{m}^{j}=f(S^{j}_{m},X^{j}_{m},\epsilon_{m}^{j})$ with $f:\mathcal{S}\times\mathcal{X}\to\mathcal{Y}$ and $\epsilon_{m}^{j}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}P_{\epsilon}$ real-valued random variables.
A widely studied special case of Assumption 2.2 considers additive, homoskedastic noise: namely, $\mathcal{Y}\subset\mathbb{R}$ and $Y_{m}^{j}=f(S^{j}_{m},X^{j}_{m})+\epsilon_{m}^{j}$. Assumption 2.2 is more general; for example, it allows the noise to be scaled by a continuous, deterministic function of $\mathcal{S}$: $Y_{m}^{j}=f(S^{j}_{m},X^{j}_{m})+g(S^{j}_{m})\epsilon_{m}^{j}$.
@@
$Q^{\textup{{test}}}\coloneqq(1/M^{\textup{{test}}})\sum\nolimits_{m=1}^{M^{\textup{{test}}}}\delta_{S^{\textup{{test}}}_{m}},$ (1)
where for $S\in\mathcal{S}$, $\delta_{S}$ denotes a Dirac measure at $S$. In this case, we call the resulting risk, written out in Def. 2.3 next, the test risk of $h$.
Definition 2.3.
Give a collection of test points $(S^{\textup{{test}}}_{m})_{m=1}^{M^{\textup{{test}}}}$ and a predictive method $h$, we define the test risk of $h$ by
$R^{\chi}_{Q^{\textup{{test}}}}(h)=(1/M^{\textup{{test}}})\sum\nolimits_{m=1}^{M^{\textup{{test}}}}e^{\chi}_{h}(S^{\textup{{test}}}_{m}).$ (2)
@@ 2.2 Estimating test risk
To estimate test risk, we assume we have access to $N^{\textup{{val}}}$ validation data points, collected in $D^{\textup{{val}}}=(S^{\textup{{val}}}_{n},X^{\textup{{val}}}_{n},Y^{\textup{{val}}}_{n})_{n=1}^{N^{\textup{{val}}}}$. We assume practitioners did not make use of either the validation or test data when constructing the predictive method.
As an example, the common holdout estimator of test risk uses the empirical average of the validation risk:
$\hat{R}^{\chi}_{\textup{{Hold}}}(h)\coloneqq(1/N^{\textup{{val}}})\sum\nolimits_{n=1}^{N^{\textup{{val}}}}\ell(Y^{\textup{{val}}}_{n},h^{\chi}(S^{\textup{{val}}}_{n})).$ (3)
@@
There exists $L>0$ such that $\forall S,S^{\prime}\in\mathcal{S}$, $|e^{\chi}_{h}(S)-e^{\chi}_{h}(S^{\prime})|\leq Ld_{\mathcal{S}}(S,S^{\prime})$.
Assumption 2.5 arises naturally in many spatial tasks. For example, consider $\mathcal{Y}\subset[0,1]$, $\mathcal{S}=(\mathbb{R}^{d},\|\cdot\|_{2})$, and squared loss. Suppose we are in the homoskedastic, additive noise setting. Suppose $\mathbb{E}[f(S,\chi(S))+\epsilon|S]$ is $L_{Y}$-Lipschitz and $h^{\chi}$ is $L_{h}$-Lipschitz. Then Assumption 2.5 holds with $L=2(L_{Y}+L_{h})$; see Prop. B.1.
@@ 3 We Want Consistent Estimators
To that end, recall the fill distance as a measure of discrepancy between two sets of points. [^3] The fill distance is not a distance in the mathematical sense since it is asymmetric and can equal 00 in cases when its two arguments are not exactly equal.
Definition 3.1 (Fill distance, Cressie 2015, §5.8, Wendland 2004, Definition 1.4).
Let $(\mathcal{S},d_{\mathcal{S}})$ be a metric space and $\Psi_{1},\Psi_{2}\subset\mathcal{S}$. The fill distance of $\Psi_{1}$ in $\Psi_{2}$ is,
$\zeta(\Psi_{1};\Psi_{2}):=\sup\nolimits_{S_{2}\in\Psi_{2}}\inf\nolimits_{S_{1}\in\Psi_{1}}d_{\mathcal{S}}(S_{1},S_{2}).$ (4)
That is, the fill distance of $\Psi_{1}$ in $\Psi_{2}$ reflects the maximum distance from a point in $\Psi_{2}$ to the nearest point in $\Psi_{1}$. [^4] The fill distance is often defined with $\Psi_{2}=\mathcal{S}$. We will sometimes require the more general definition.
In the spatial statistics literature, infill asymptotics describes cases where data are gathered over a compact spatial domain in such a way that the fill distance of the data to its domain tends to 00 (Cressie, 2015, Chapter 5.8). We will say an estimator is consistent for the test risk under infill asymptotics if – for any $Q^{\textup{{test}}}$, $\chi$, and $h$ satisfying our assumptions above – the estimator converges in probability to $R^{\chi}_{Q^{\textup{{test}}}}(h)$.
@@
Fix a predictive method $h$ and a test measure $Q^{\textup{{test}}}$. Take Assumptions 2.2, 2.1, 2.4 and 2.5. Consider an infinite sequence of validation sets of increasing size: $(D^{\textup{{val}}}_{N})_{N=1}^{\infty},D^{\textup{{val}}}_{N}=(S^{\textup{{val}}}_{N},X^{\textup{{val}}}_{N},Y^{\textup{{val}}}_{N})_{n=1}^{N}$ such that when $N^{\prime}<N$, the first $N^{\prime}$ points of $D^{\textup{{val}}}_{N}$ are $D^{\textup{{val}}}_{N^{\prime}}$. Suppose $\lim_{N\to\infty}\zeta(S^{\textup{{val}}}_{1:N},\mathcal{S})=0$. Let $\hat{R}_{N}$ be an estimator constructed from the validation data $D^{\textup{{val}}}_{N}$. We say that the estimator $\hat{R}_{N}$ is consistent for the test risk under infill asymptotics if $\lim_{N\to\infty}\mathrm{Pr}(|\hat{R}_{N}-R^{\chi}_{Q^{\textup{{test}}}}(h)|\geq\epsilon)=0.$
That is, as validation data fills the spatial domain, the estimator should converge to the test risk – no matter the composition of test sites. Note that we do not require convergence to be uniform over test measures. Also note that our assumption that fill distance tends to zero is generally weaker than an assumption that the validation sites are drawn i.i.d. from a distribution with Lebesgue density supported on the spatial domain. Reznikov & Saff (2015, Theorem 2.1) showed an implication relationship between these assumptions in a much more general setting for much more general metric spaces, and Vacher et al. (2021, Lemma 12) discusses the special case for the unit cube. Next, we present a finite sample version of this implication, with an advantage relative to past work that we keep track of all constants.
@@
Suppose that $\mathcal{S}=[0,1]^{d}$, $S^{\textup{{val}}}_{n}\stackrel{{\scriptstyle\textup{{iid}}}}{{\sim}}P$ for $1\leq n\leq N^{\textup{{val}}}$, and $P$ has Lebesgue density lower bounded by $c>0$ over $[0,1]^{d}$. Let $B_{d}=\pi^{d/2}/\Gamma(\frac{d}{2}+1)$ be the volume of the $d$-dimensional Euclidean unit ball. For any $\delta\in(0,1)$ there exists an $n_{0}$ such that for all $N^{\textup{{val}}}\geq n_{0}$ with probability at least $1-\delta$
$\zeta(S^{\textup{{val}}}_{1:N^{\textup{{val}}}},[0,1]^{d})\leq\left(\tfrac{4^{d}}{cN^{\textup{{val}}}B_{d}}\left(\log\tfrac{6^{d}N^{\textup{{val}}}}{B_{d}\delta}\right)\right)^{1/d}.$ (5)
We prove Prop. 3.3 in Section B.2. The right-hand side of this bound is $O((\log N^{\textup{{val}}}/N^{\textup{{val}}})^{1/d})$, and so the fill distance converges to zero in probability under these assumptions.
Consistency under infill asymptotics is a minimal desirable property.
Much like traditional consistency, we emphasize that consistency as described here is just a single check among many. For instance, often practitioners will be interested in extrapolation far from observed data, which is not modelled by infill asymptotics and will need to be considered separately. Our only supposition here is that we will generally prefer test-risk estimators that satisfy consistency under infill asymptotics to those that do not.

### v2

@@ Abstract
Spatial prediction tasks are key to weather forecasting, studying air pollution impacts, and other scientific endeavors. Determining how much to trust predictions made by statistical or physical methods is essential for the credibility of scientific conclusions. Unfortunately, classical approaches for validation fail to handle mismatch between locations available for validation and (test) locations where we want to make predictions. This mismatch is often not an instance of covariate shift (as commonly formalized) because the validation and test locations are fixed (e.g., on a grid or at select points) rather than i.i.d. from two distributions. In the present work, we formalize a check on validation methods: that they become arbitrarily accurate as validation data becomes arbitrarily dense. We show that classical and covariate-shift methods can fail this check. We instead propose a method that builds from existing ideas in the covariate-shift literature, but adapts them to the validation data at hand. We prove that our proposal passes our check. And we demonstrate its advantages empirically on simulated and real data.
[…]
@@ 1 Introduction
Researchers are often interested in making predictions in a spatial setting. For instance, scientists predict sea surface temperature (SST) for weather forecasting and climate research (Minnett, 2010), predict air pollution at population centers to better understand the effect of pollution on health outcomes such as kidney disease (Remigio et al., 2022), or predict the prevalence of an invasive species for ecological management (Barbet-Massin et al., 2018). Characterizing the reliability of these predictions is key to understanding their suitability for downstream applications; e.g., Minnett (2010) describes acceptable SST error tolerances for weather forecasting. Estimates of prediction accuracy can also be used to choose between several predictive methods, as in Shabani et al. (2016).
@@
In the spatial setting, predictive methods need not always arise from a statistical or machine learning approach built using training data. The predictive method is often a complex physical model provided by a third party (Remigio et al., 2022; Minnett, 2010; Gupta et al., 2018). Or it could combine physics and data-driven models (Banzon et al., 2016; Werner et al., 2019; Özkaynak et al., 2013).
@@
In any of these cases, it is common to estimate the performance of a predictive method by using a set of validation data. More precisely, we are ultimately interested in predicting a response at what we call test sites; in the SST example above (Minnett, 2010), the test sites are points on a grid (often called a map), or in the air pollution example (Remigio et al., 2022), the test sites are 28 counties in the US Northeast. We can make predictions at the test sites, but we do not have access to direct observations of the responses there. We do have observed responses in the validation data, which we assume were not used in forming the predictive method being evaluated; in the SST example, scientists have SST observations taken by boats and buoys as validation data. So our aim is to estimate the average loss (i.e. risk) at the test sites using the validation data. We will see that many popular or natural approaches fail at this task.
@@
One widely used approach, called the holdout, [^1] The name originally referred to “holding out” data for validation, with the remainder of available data going toward training. While we maintain the naming convention, we emphasize that in our setup there need not be any training data. Minnett (2010); Gupta et al. (2018); Duan et al. (2019), among many others, use this approach. estimates the test risk by taking the empirical average of the validation loss. When the validation and test data are independent and identically distributed (i.i.d.) from the same distribution, the holdout has a rigorous justification (Devroye, 1976; Langford, 2005). But in spatial problems, the validation and test sites need not be similarly dispersed, all data may be spatially correlated, and the test sites are often fixed rather than random; recall the grid or point prediction examples above. Indeed, Roberts et al. (2017) have observed problems with the holdout in practice. In the special case where the predictive method is data-driven, some authors (e.g. Telford and Birks, 2005) have suggested choosing holdout validation sites far from the training data sites. But we expect this proposal to still suffer from the problems just mentioned, and in fact simulation studies suggest it can misestimate test risk (Ploton et al., 2020; De Bruin et al., 2022).
@@
In what follows, we lay out a precise formulation of the prediction validation task in the spatial setting (Section 2). We formalize a desirable property for test-risk estimators: that, if arbitrarily dense validation data accrues in a region including the test points, the test-risk estimate should become arbitrarily accurate (Section 3). We prove that both the holdout estimator and an estimator advocated in the covariate-shift literature (Loog, 2012; Portier et al., 2023) fail to satisfy this spatial consistency property (Section 4). We propose to build on the $k$-nearest neighbor estimator (Loog, 2012). In particular, Loog (2012) and Portier et al. (2023) advocated fixing $k=1$ for covariate-shift problems. We instead derive an upper bound on the error of the general-$k$ estimator for estimating test risk (Section 5.1); crucially, our bound is conditional on the test and validation sites. We prove that choosing $k$ adaptively by optimizing our upper bound yields a spatially consistent estimator (Section 5.2). Unlike covariate-shift results (e.g. Portier et al., 2023), our results are directly applicable to problems where the test sites are most reasonably thought of as fixed. We illustrate the accuracy and practicality of our proposed method in simulated and real data analyses (Section 6), with tasks in both grid and point prediction. We discuss further related work in Appendix B.
@@ 2 Estimating Test Risk in a Spatial Problem
We now formalize risk estimation at test points in a spatial setting. We assume each data point occurs at a spatial location $S\in\mathcal{S}$, where the spatial domain $(\mathcal{S},d_{\mathcal{S}})$ is a metric space. Each data point has observed covariates $X\in\mathcal{X}$ and a response $Y\in\mathcal{Y}$. The covariates are a fixed spatial field, $\chi:\mathcal{S}\to\mathcal{X}$; i.e., the covariates at a point are specified by evaluating $\chi$ at the point’s spatial location.
@@ 2.1 Test risk of a spatial predictive method
We assume we have access to a predictive method $h:(\mathcal{S},\mathcal{X})\to\mathcal{Y}$. We define $h^{\chi}:\mathcal{S}\to\mathcal{Y}$ to be the prediction made by $h$ at the location $S$: $h^{\chi}(S)=h(S,\chi(S))$. We suppose that practitioners would like to use $h$ to predict the response at a set of test sites where the response is unknown. We collect the $M^{\textup{{test}}}$ test data points, including true (but unobserved) responses, in $\smash{D^{\textup{{test}}}=(S^{\textup{{test}}}_{m},X^{\textup{{test}}}_{m},Y^{\textup{{test}}}_{m})_{m=1}^{M^{\textup{{test}}}}}$.
To quantify quality of a predictive method, we need a loss. We assume the loss is bounded, as is often the case for practically-bounded responses; cf. temperature, pressure, or other physical quantities. [^2] Loss is also bounded for classification error or robust regression cases such as Tukey’s biweight loss.
@@
The loss is a non-negative, bounded function, $\ell\!:\!\mathcal{Y}\times\!\mathcal{Y}\to\![0,\Delta]$.
@@
Let $j=\text{test}$. Let $\chi:\mathcal{S}\to\mathcal{X}$ be a fixed function. For $1\leq m\leq M^{j}$, $X^{j}_{m}=\chi(S^{j}_{m})$ and $Y_{m}^{j}=f(S^{j}_{m},X^{j}_{m},\epsilon_{m}^{j})$ with $f:\mathcal{S}\times\mathcal{X}\to\mathcal{Y}$ and $\epsilon_{m}^{j}\smash{\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}}P_{\epsilon}$ real-valued random variables.
Assumption 2.2 implies the response is i.i.d. given the spatial location. For example, Assumption 2.2 would cover the case where measurement errors on sensors are independent, but the locations of the sensors are not. A widely studied special case of Assumption 2.2 considers additive, homoskedastic noise: namely, $\mathcal{Y}\subset\mathbb{R}$ and $Y_{m}^{j}=f(S^{j}_{m},X^{j}_{m})+\epsilon_{m}^{j}$. Assumption 2.2 is more general; for example, it allows the noise to be scaled by a continuous, deterministic function of $\mathcal{S}$: $Y_{m}^{j}=f(S^{j}_{m},X^{j}_{m})+g(S^{j}_{m})\epsilon_{m}^{j}$. Before defining risk, we first define the average loss of the predictive method at a particular location in space, $S$: $e_{h}(S):=\mathbb{E}[\ell(f(S,\chi(S),\epsilon),h^{\chi}(S))|S]$. Finally, we average over all spatial locations of interest, which we assume is a finite set.
Definition 2.3.
@@ 2.2 Estimating test risk
To estimate test risk, we assume we have access to $N^{\textup{{val}}}$ validation data points, collected in $\smash{D^{\textup{{val}}}=(S^{\textup{{val}}}_{n},X^{\textup{{val}}}_{n},Y^{\textup{{val}}}_{n})_{n=1}^{N^{\textup{{val}}}}}$. We assume practitioners did not make use of either the validation or test response data when constructing the predictive method. As an example, the common holdout estimator uses the empirical average of validation loss:
$\smash{\hat{R}_{\textup{{Hold}}}(h)\coloneqq(1/N^{\textup{{val}}})\sum\nolimits_{n=1}^{N^{\textup{{val}}}}\ell(Y^{\textup{{val}}}_{n},h^{\chi}(S^{\textup{{val}}}_{n})).}$ (1)
@@
For some $L\geq 0$, for all $S,S^{\prime}\in\mathcal{S},|e_{h}(S)\!-\!e_{h}(S^{\prime})|\leq Ld_{\mathcal{S}}(S,S^{\prime}).$
Assumption 2.5 often arises naturally. For example, consider $\mathcal{Y}\subset[0,1]$, $\mathcal{S}=(\mathbb{R}^{d},\|\cdot\|_{2})$, squared loss, and homoskedastic and additive noise. Suppose $f(S,\chi(S))$ is $L_{Y}$-Lipschitz and $h^{\chi}$ is $L_{h}$-Lipschitz. Then Assumption 2.5 holds with $L=2(L_{Y}+L_{h})$; see Proposition D.1.
@@ 3 We Want Consistent Estimators
To that end, note that the fill distance is a measure of discrepancy between two sets of points, $\Psi_{1}$ and $\Psi_{2}$. [^3] The fill distance is not a distance in the mathematical sense since it is asymmetric and can equal 00 in cases when its two arguments are not exactly equal. It is the maximum distance from a point in $\Psi_{2}$ to the nearest point in $\Psi_{1}$.
Definition 3.1 (Cressie 2015, §5.8, Wendland 2004, Definition 1.4).
Let $(\mathcal{S},d_{\mathcal{S}})$ be a metric space and $\Psi_{1},\Psi_{2}\subset\mathcal{S}$. The fill distance of $\Psi_{1}$ in $\Psi_{2}$ is $\zeta(\Psi_{1};\Psi_{2}):=\sup\nolimits_{S_{2}\in\Psi_{2}}\inf\nolimits_{S_{1}\in\Psi_{1}}d_{\mathcal{S}}(S_{1},S_{2}).$
In the spatial statistics literature, infill asymptotics describes cases where data are gathered over a compact spatial domain in such a way that the fill distance of the data to its domain tends to 00 (Cressie, 2015, §5.8). We say an estimator is consistent for the test risk under infill asymptotics if – for any $Q^{\textup{{test}}}$, $\chi$, and $h$ satisfying our assumptions above – the estimator converges in probability to $R_{Q^{\textup{{test}}}}(h)$.
@@
Fix a predictive method $h$ and a test measure $Q^{\textup{{test}}}$. Take Assumptions 2.2, 2.1, 2.4 and 2.5. Consider an infinite sequence of validation sets of increasing size: $(D^{\textup{{val}}}_{N})_{N=1}^{\infty},D^{\textup{{val}}}_{N}=(S^{\textup{{val}}}_{n},X^{\textup{{val}}}_{n},Y^{\textup{{val}}}_{n})_{n=1}^{N}$ such that when $N^{\prime}<N$, the first $N^{\prime}$ points of $D^{\textup{{val}}}_{N}$ are $D^{\textup{{val}}}_{N^{\prime}}$. Suppose $\smash{\lim_{N\to\infty}\zeta(S^{\textup{{val}}}_{1:N},\mathcal{S})=0}$. Let $\smash{\hat{R}_{N}}$ be an estimator constructed from the validation data $D^{\textup{{val}}}_{N}$. We say that the estimator $\smash{\hat{R}_{N}}$ is consistent for the test risk under infill asymptotics if for all $\epsilon>0,\lim_{N\to\infty}\mathrm{Pr}(|\smash{\hat{R}_{N}}-R_{Q^{\textup{{test}}}}(h)|\geq\epsilon)=0.$
That is, as validation data fills the spatial domain, the estimator should converge to the test risk – no matter the composition of test sites. Our assumption that fill distance tends to zero is generally weaker than an assumption that the validation sites are drawn i.i.d. from a distribution with Lebesgue density supported on the spatial domain. Reznikov and Saff (2015, Theorem 2.1) showed an implication relationship between these assumptions in a much more general setting, and Vacher et al. (2021, Lemma 12) discuss the special case for the unit cube. Next, we present a finite-sample version of this implication, with an advantage relative to past work that we keep track of all constants.
@@
Suppose that $\mathcal{S}=[0,1]^{d}$, $S^{\textup{{val}}}_{n}\smash{\stackrel{{\scriptstyle\textup{{\tiny iid}}}}{{\sim}}}P$ for $1\leq n\leq N^{\textup{{val}}}$, and $P$ has Lebesgue density lower bounded by $c>0$ over $\smash{[0,1]^{d}}$. Let $B_{d}=\smash{\pi^{d/2}}/\Gamma(d/2+1)$ be the volume of the $d$-dimensional Euclidean unit ball. For any $\delta\in(0,1)$ there exists an $n_{0}$ such that for all $N^{\textup{{val}}}\geq n_{0}$ with probability at least $1-\delta$
$\smash{\zeta(S^{\textup{{val}}}_{1:N^{\textup{{val}}}},[0,1]^{d})\leq\left(\tfrac{4^{d}}{cN^{\textup{{val}}}B_{d}}\left(\log\tfrac{6^{d}N^{\textup{{val}}}}{B_{d}\delta}\right)\right)^{1/d}.}$ (2)
We prove Proposition 3.3 in Section D.2. The right side of this bound is $\smash{O((\log N^{\textup{{val}}}/N^{\textup{{val}}})^{1/d})}$, and so the fill distance converges to zero in probability under these assumptions.
Consistency under infill asymptotics is a minimal desirable property. Like traditional consistency, we emphasize that Definition 3.2 is just a single check among many. For instance, often practitioners will be interested in extrapolation far from observed data, which is not modelled by infill asymptotics and will need to be considered separately. Our only supposition here is that we will generally prefer test-risk estimators that satisfy consistency under infill asymptotics to those that do not.